# Leveraging Data Augmentation and Siamese Learning for Predictive Process Monitoring

### v1

@@ 6.0.6 Ablation Study
Figure 7 illustrates the impact of data augmentation on model accuracy using varying fractions (20%, 40%, 60%) of the BPIC15-1 training data. We compare a Transformer model and a Random Forest classifier, each trained on raw subsets, subsets augmented using a random strategy and subsets augmented using our proposed method. For the Transformer, our method consistently improves accuracy at 20% and 40% sample sizes, bringing performance close to that achieved with the full 80% training data. Interestingly, on the smallest subset (20%), random augmentation performs slightly better, likely due to the limited information in the fewer number of traces. In contrast, the Random Forest classifier shows no performance gains with either augmentation method, remaining well below the full-data baseline. These results suggest that deep learning models are better equipped to exploit synthetic variation introduced through augmentation, leading to better performance.

### v2

@@ 6.0.6 Ablation Study
Fig.  7 illustrates the impact of data augmentation on model accuracy using varying fractions (20%, 40%, 60%) of the BPIC15-1 training data. We compare a Transformer model and a Random Forest classifier, each trained on raw subsets, subsets augmented using a random strategy and subsets augmented using our proposed method. For the Transformer, our method consistently improves accuracy at 20% and 40% sample sizes, bringing performance close to that achieved with the full 80% training data. Interestingly, on the smallest subset (20%), random augmentation performs slightly better, likely due to the limited information in the fewer number of traces. In contrast, the Random Forest classifier shows no performance gains with either augmentation method, remaining well below the full-data baseline. These results suggest that deep learning models are better equipped to exploit synthetic variation introduced through augmentation, leading to better performance.